Fermi National Laboratory


Accelerator Update for January 10, 2005 - January 12, 2005

Monday January 10
The day shift began with Operations monitoring store 3917, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

The Recycler began stashing antiprotons at 9:43 AM.

The stashing ended and Pbar resumed stacking at 10:24 AM.

Operations established beam to the Meson MTest experiment at 11:13 AM.

A NuMI expert began a study period at 1:26 PM.

Operations established beam to the Meson MCenter experiment at 6:31 PM.

At 11:12 PM, the air compressor at the Tevatron (TeV) F3 service building alarmed. The Duty Electrician investigated, but found nothing wrong and the alarm cleared.

Tuesday January 11
The midnight shift began with Operations monitoring store 3917, with Pbar stacking, and with MiniBooNE taking beam.

MTest experimenters finished using beam at 12:02 AM.

At 1:28 AM, the F3 air compressor alarmed again. Operator investigated and observed the compressor operating oddly for a few cycles and then resumed normal operation. This information was passed on and, later, technicians fixed the compressor's drier.

At 2:04 AM, a Recycler power supply (R:CE304) tripped off. An expert came in and changed out a control chassis. By the time this was finished at 2:25 AM, the Recycler had lost 27E10 from its stash.

Operations terminated store 3917 at 7 AM.

Cryo system technicians completed their repairs on the TeV sector C1 wet engine at 8:02 AM.

An expert reported at 8:47 AM that the F3 air compressor was repaired.

Operations put the TeV into shot setup around 8 AM and then established store 3918 at 10:19 AM. Its initial luminosity was 69.88E30.

At 10:28 AM, a Main Injector (MI) RF station (MIRF2) tripped off due to a bad modulator. MI continued with 16 RF stations.

Operations established beam to MTest at 10:49 AM.

At 11:30 AM, Operations reported that a MI low level RF frequency problem caused all the MI RF stations to trip off. Operators reset the system.

At 12:30 PM, the MI frequency problem again tripped off the RF stations. This happened twice more and then the system wouldn't reset. An expert shut down a process running on a PC. He commented later that all the stations were running harder because of the two stations that were down.

Pbar experts announced at 12:30 PM, that Pbar achieved a new stacking record: 13.78mA per hour.

At 1:13 PM, the MI RF was shut off to repair the MIRF2 modulator. The station was back on twenty minutes later.

Pbar began a reverse proton study at 4:16 PM. This study halted stacking. Pbar resumed stacking at 10:22 PM.

Operations put the Recycler into shot setup at 11:33 PM.

Wednesday January 12
The midnight shift began with Operations monitoring store 3918, and with the Recycler stashing antiprotons.

Stashing ended and stacking resumed at 12:50 AM.

Operations established beam to MCenter at 1 AM.

At 5:31 AM, Operations reported that Pbar's Debuncher RF stations 1-6 were off. Operators tried various things to fix a driver, but ended up with all the original equipment back in place and the driver working normally. (They passed this problem on to the expert in the morning, who fixed it by adjusting the current.)

Operations reported at 7:34 AM that the Neutron Therapy Facility (NTF) was setting up to take beam.

The Plan for Wednesday
The plan for today is to terminate the present store at 5 PM and then put the TeV into a mixed antiproton shot setup.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/12/2005   email Fermilab